# Optimal Control of an SIR Model with Noncompliance as a Social Contagion

### v1

@@ 2. An SIR Model with Behavioral Effects: Disease-Free Equilibria and Reproductive Ratios
Remark. This proposition is a sanity check to verify that our model behaves as expected. The two points can be interpreted as follows: (i) more control leads to a lower rate of disease spread, and (ii) when we are at the equilibrium population level of $\tfrac{b}{\delta}$, more noncompliance leads to a higher rate of disease spread.
Proof.
That $\mathscr{R}_{0}(s,s^{*})$ is decreasing in $\alpha$ when other values are held constant is obvious directly from (2.11) since $\mathscr{R}_{0}(s,s^{*})$ is linear in $\alpha$. That $\mathscr{R}_{0}(s,s^{*})$ is decreasing in $\mu$ can be seen from the second line (2.11) since increasing $\mu$ increases each of the nested fractions, each of which has the effect of decreasing $\mathscr{R}_{0}(s,s^{*}).$ The remaining statements in (i), (ii) are not so obvious directly from the formula, but follow from the observation that $z\mapsto\frac{c_{1}+c_{2}z}{c_{3}+c_{4}z}$ is increasing for $z\geq 0$ when $c_{2}c_{3}\geq c_{1}c_{4}$ and decreasing for $z\geq 0$ when $c_{2}c_{3}\leq c_{1}c_{4}$. For (i), alternately writing $\mathscr{R}(s,s^{*})$ in this form where the role of $z$ is played by $\eta$ or $\nu$, the inequalities proving that $\mathscr{R}_{0}(s,s^{*})$ is decreasing boils down to
$(\gamma+\delta)(\gamma+\nu+\delta+(\overline{\mu}-\mu)s^{*})s^{*}$ $\leq(\gamma+\delta+\nu)(\gamma+\nu+\delta+(\overline{\mu}-\mu)s^{*})((1-\alpha)s+s^{*}),$
$(\gamma+\delta)(\gamma+\eta+\delta+(\overline{\mu}-\mu)s^{*})$ $\leq(\gamma+\eta+\delta)\Big{(}\gamma+\delta+(\overline{\mu}-\mu)s^{*}+\tfrac{\eta s^{*}}{(1-\alpha)s+s^{*}}\Big{)},$
respectively. Each of these is verified simply by checking that the right-hand side has all the same terms as the left hand side plus a few extra positive terms.
For part (ii), if $s^{*}\in[0,\tfrac{b}{\delta}]$, we see
(2.12) $\mathscr{R}_{0}(\tfrac{b}{\delta}-s^{*},s^{*})=\frac{\beta\Big{[}\tfrac{b}{\delta}(\gamma+\nu+\delta)+\Big{(}\tfrac{b}{\delta}(\overline{\mu}-\mu)+\eta\Big{)}s^{*}\Big{]}}{\Big{(}(\gamma+\delta)(\gamma+\eta+\nu+\delta)+\eta\nu\Big{)}+(\gamma+\delta)(\overline{\mu}-\mu)s^{*}}.$
That this is increasing in $s^{*}$ follows from the same reasoning with the inequality flipped. ∎
To conclude this section, we state stability results for our DFE. We state the theorem here, but relegate the proof to an appendix since it requires a long computation.
@@
    If $\xi=0$ and $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$, then $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$ and unstable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)>1$.
@@
    If $\xi=0$ and $\tfrac{b}{\delta}>\tfrac{\nu+\delta}{\overline{\mu}-\mu}$, then $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})<1$ and unstable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})>1$
@@
Remark. In each case, that the condition on the reproductive ratio determines local asymptotic stability or instability follows directly from [47, Theorem 2] once the assumptions of that theorem are verified. However, the stability is actually global, in the sense that under the conditions of each statement, any solution of (2.2) will tend toward the given equilibrium point regardless of whether the initial condition is nearby the equilibrium point. This global version can be proven using appropriately constructed Lyapunov functions as demonstrated in the appendix and in [39].
@@ 3. Optimal Control of our Model
With this notation, we define the marginal cost function,
(3.1) $r(x,u)=c_{1}(I+I^{*})+c_{2}N^{*}+\frac{c_{3}}{2}\alpha^{2}+\frac{c_{4}}{2}\eta^{2}+\frac{c_{5}}{2}\mu^{2}+\frac{c_{6}}{2}\nu^{2},$
where, as above, $N^{*}=S^{*}+I^{*}+R^{*}$. Our cost functional is then
(3.2) $J(x,u)=\int^{T}_{0}r(x(t),u(t))dt.$
In (3.1), $c_{1},\ldots,c_{6}$ are positive weights indicating the preferences of the policy-maker. For example, a more public-health oriented policy-maker would take $c_{1}$ larger, whereas an economically stingy policy-maker would increase $c_{3},c_{4},c_{5},c_{6}$. From the above, our admissible control set is
(3.3) $U_{\text{ad}}=\{(\alpha,\eta,\mu,\nu)\in{\mathbb{R}}^{4}\,:\,0\leq\alpha\leq 1,\,0\leq\eta\leq\overline{\eta},\,0\leq\mu\leq\overline{\mu},\,0\leq\nu\leq\overline{\nu}\}$
@@
(3.4) $\mathscr{U}_{\text{ad}}=\{u(\cdot)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(\alpha(\cdot),\eta(\cdot),\mu(\cdot),\nu(\cdot)):[0,T]\to U_{\text{ad}}\,\,|\,\,u\text{ measurable}\}.$
Our optimal control problem is then
@@
where $f(x,u)$ is defined by the right-hand side of (2.2) and $x_{0}\in{\mathbb{R}}^{6}$ has nonnegative components whose sum is bounded by $\tfrac{b}{\delta}$.
Existence of an optimal control for this problem follows from classical optimal control theory.
@@
Assuming $c_{3},c_{4},c_{5},c_{6}$ are positive and $x_{0}$ has nonnegative components, there exists an optimal control $u\in\mathscr{U}_{\text{ad}}$ which solves the minimization problem (3.5).
We omit the proof. Some key features of our system are that solutions of our system remain bounded and nonnegative as long as $x_{0}$ is componentwise nonnegative, $f(x,u)$ is smooth and affine in the control variables, and $r(x,u)$ is smooth and uniformly strongly convex in $u$ (when each of $c_{3},c_{4},c_{5},c_{6}$ are positive), and the control set $U_{\text{ad}}$ is convex and compact. These suffice to prove existence of an optimal control using [45, Theorem 2.1]. In fact the particular dependence of running cost (3.1) on the state variables $x=(S,I,R,S^{*},I^{*},R^{*})$ could be significantly altered, as long as the smoothness in all variables and uniform strong convexity in control variables are maintained. We choose this particular running cost only to represent the concerns we imagine a policy-maker may have; other decisions are certainly valid.
@@
A necessary condition for optimality is provided by the Pontryagin minimum principle (PMP). Using state variables $x=(S,I,R,S^{*},I^{*},R^{*})$, costate variables $p=(p_{S},p_{I},p_{R},p_{S^{*}},p_{I^{*}},p_{R^{*}})$, and control variables $u=(\alpha,\eta,\mu,\nu),$ we define the Hamiltonian
(3.6) $\begin{split}H(x,p,u)&=\langle p,f(x,u)\rangle+r(x,u),\end{split}$
@@
where $r(x,u)$ is the running cost (3.1) and $f(x,u)$ is defined by the right-hand side of (2.2). Then the PMP states that if $x(\cdot),u(\cdot)$ are an optimal state and control pair for (3.5), then there is a corresponding costate trajectory $p(\cdot)$ such that

### v2

@@ 2. An SIR Model with Behavioral Effects: Disease-Free Equilibria and Reproductive Ratios
Remark. This proposition is a sanity check to verify that our model behaves as expected. The two points can be interpreted as follows: (i) more control leads to a lower rate of disease spread, and (ii) when we are at the equilibrium population level of $\tfrac{b}{\delta}$, more noncompliance leads to a higher rate of disease spread. We omit the proof because it is elementary, if tedious. It follows most readily from the observation that, the function $z\mapsto\frac{c_{1}+c_{2}z}{c_{3}+c_{4}z}$ in increasing for $z\geq 0$ when $c_{2}c_{3}\geq c_{1}c_{4}$, and decreasing for $z\geq 0$ when $c_{2}c_{3}\leq c_{1}c_{4}$. Clearing the nested denominators, one can write $\mathscr{R}_{0}(s,s^{*})$ in this form with $\alpha,\eta,\mu,\nu$ alternately playing the role of $z$, and verify the correct inequality.
To conclude this section, we state stability results for our DFE.
@@
    If $\xi=0$ and $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\mu^{*}-\mu}$, then $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$ and unstable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)>1$.
@@
    If $\xi=0$ and $\tfrac{b}{\delta}>\tfrac{\nu+\delta}{\mu^{*}-\mu}$, then $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})<1$ and unstable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})>1$
@@
Remark. In each case, that the condition on the reproductive ratio determines local asymptotic stability or instability follows directly from [55, Theorem 2] once the assumptions of that theorem are verified. However, the stability is actually global, in the sense that under the conditions of each statement, any solution of (2.2) will tend toward the given equilibrium point regardless of whether the initial condition is nearby the equilibrium point. This global version can be proven using appropriately constructed Lyapunov functions. We omit the proof because it is quite long and technical, and is contained in [45].
@@ 3. Optimal Control of our Model
With this notation, we define a generic cost functional
(3.1) $J(x,u)=\int^{T}_{0}(r(x(t))+\ell(u(t)))dt$
where $r:{\mathbb{R}}^{6}_{\geq 0}\to{\mathbb{R}}$ and $\ell:{\mathbb{R}}^{4}\to{\mathbb{R}}$. We state assumptions on $r,\ell$ in theorem 3.1, but we make special note that we do not require strict or strong convexity. The exact choice of $r,\ell$ indicate the preferences of the policy-maker. For example, a more public-health oriented policy-maker would choose $r$ which assigns large values to states with higher $I$ and $I^{*}$, whereas an economically stingy policy-maker would choose $\ell$ to take large values for larger values of the controls. From the above, our admissible control set is
(3.2) $U_{\text{ad}}=\{(\alpha,\eta,\mu,\nu)\in{\mathbb{R}}^{4}\,:\,0\leq\alpha\leq\overline{\alpha},\,0\leq\eta\leq\overline{\eta},\,0\leq\mu\leq\overline{\mu},\,0\leq\nu\leq\overline{\nu}\}$
@@
(3.3) $\mathscr{U}_{\text{ad}}=\{u(\cdot)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(\alpha(\cdot),\eta(\cdot),\mu(\cdot),\nu(\cdot))\in BV([0,T];U_{\text{ad}})\,:\,TV(\alpha),TV(\eta),TV(\nu),TV(\mu)\leq M\}$
for some universal constant $M$. That is, each individual control $\alpha,\eta,\mu,\nu$ is in $BV[0,T]$, the space of real-valued functions of bounded variation. To recall, this is the set of functions $z:[0,T]\to{\mathbb{R}}$ such that
(3.4) $TV(z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\sup_{g\in C_{c}^{1}[0,T],\|g\|_{\infty}\leq 1}\int^{T}_{0}z(t)g^{\prime}(t)dt<\infty.$
We remark on this choice of control functions after the proof of theorem 3.1. Our optimal control problem is then
@@
where $f(x,u)$ is defined by the right-hand side of (2.2) and $x_{0}\in{\mathbb{R}}^{6}$ has nonnegative components whose sum is bounded by $\tfrac{b}{\delta}$. In this notation, we note that the unique a.e. solution $x$ of (2.2) corresponding to a control map $u\in\mathscr{U}_{ad}$ can be characterized as the unique function $x\in(L^{2}[0,T])^{6}$ satisfying the initial condition and
(3.6) $\int^{T}_{0}\langle x(t),\phi^{\prime}(t)\rangle dt=\int^{T}_{0}\langle f(x(t),u(t)),\phi(t)\rangle dt,\,\,\,\,\,\forall\phi\in C^{1}_{c}([0,T];{\mathbb{R}}^{6}).$
Because our admissible control set $U_{\text{ad}}$ is compact, $\mathscr{U}_{\text{ad}}\subset(L^{p}[0,T])^{4}$ for any $p\in[1,\infty)$ or $p=\infty.$ For simplicity, we prove the following existence theorem using the direct method considering convergence in the $L^{2}$ topology, though with small tweaks this argument would go through with other choices.
@@
Assuming $r$ and $\ell$ are proper, convex, and lower-semicontinuous and $x_{0}$ has nonnegative components, there exists an optimal control $u^{*}\in\mathscr{U}_{\text{ad}}$ and corresponding trajectory $x^{*}$ which solve the minimization problem (3.5).
Proof.
By proposition 2.1, there is a well-defined control-to-state operator $\mathcal{S}:\mathscr{U}_{\text{ad}}\to(H^{1}[0,T])^{6}$ so that for $u\in\mathscr{U}_{\text{ad}}$, $\mathcal{S}(u)=x\in(H^{1}[0,T])^{6}$ is the corresponding solution of (2.2). Note that the basic existence theory will only guarantee that $x\in(L^{2}[0,T])^{6}$, but then by the population bound (2.3), the right hand side of each equation in (2.2) defines a bounded, and hence $L^{2}[0,T]$, function so $x\in(H^{1}[0,T])^{6}$ by a bootstrapping argument. Furthermore $\mathcal{S}(u)=x$ remains uniformly bounded in $(H^{1}[0,T])^{6}$ among choices of $u\in\mathscr{U}_{\text{ad}}.$ We segment the proof into a few steps. In all following statements, we assume $\mathscr{U}_{\text{ad}}$ and $(H^{1}[0,T])^{6}$ are given the $L^{2}$ topology.
Step 1: $\mathcal{S}$ is weakly closed.
Suppose $u^{(k)}\in\mathscr{U}_{\text{ad}}$ is a sequence such that $u^{(k)}\rightharpoonup u$ in $(L^{2}[0,T])^{4}$ and $\mathcal{S}(u^{(k)})\rightharpoonup x$ in $(L^{2}[0,T])^{6}.$ Since $\mathscr{U}_{\text{ad}}$ is convex, it suffices to prove that it is strongly closed in $L^{2}$, and this will imply that it is weakly closed [15, Ch. 5.1]. That $\mathscr{U}_{\text{ad}}$ is closed under strong $L^{2}$ limits follows because if $z^{(k)}\to z$ strongly in $L^{2}$, then Hölder’s inequality gives
$\left|\int^{T}_{0}z^{(k)}(t)g^{\prime}(t)dt-\int^{T}_{0}z(t)g^{\prime}(t)dt\right|\leq\|z^{(k)}-z\|_{L^{2}[0,T]}\|g^{\prime}\|_{L^{2}[0,T]}\to 0$
for any $g\in C_{c}^{1}[0,T]$, and this proves that $TV(z^{(k)})\to TV(z)$ for total variation as defined in (3.4). In particular, strong $L^{2}$ limits respect bounds on total variation. Hence $\mathscr{U}_{\text{ad}}$ is weakly closed and so $u\in\mathscr{U}_{\text{ad}}$. Further, since $\{u^{(k)}\}$ is of uniformly bounded total variation, Helly’s selection theorem [25, Thm. 5.5] allows us to pass to a subsequence along which $u^{(k)}\to u$ pointwise a.e. in $[0,T].$ (We do so without renaming the sequence, as we do whenever we mention a subsequence in the ensuing work.)
Now $\mathcal{S}(u^{(k)})$ is a bounded sequence in the Hilbert space $(H^{1}[0,T])^{6}$, so it has subsequence converging weakly to some $\overline{x}\in(H^{1}[0,T])^{6}$. By the compact embedding $H^{1}[0,T]\hookrightarrow L^{2}[0,T]$ [24, Ch. 5.7], passing to a further subsequence, we can ensure strong convergence to $\overline{x}$ in the $(L^{2}[0,T])^{6}$ topology. By uniqueness of limits, we must have $\overline{x}=x$, and thus passing to an even further subsequence, we have $\mathcal{S}(u^{(k)})\to x$ pointwise a.e.
Finally, for each $k$, using the weak formulation (3.6), we have
(3.7) $\int^{T}_{0}\langle[\mathcal{S}(u^{(k)})](t),\phi^{\prime}(t)\rangle dt=\int^{T}_{0}\langle f([\mathcal{S}(u^{(k)})](t),u^{(k)}(t)),\phi(t)\rangle dt,\,\,\,\,\,\forall\phi\in C^{1}_{c}([0,T];{\mathbb{R}}^{6}).$
As $k\to\infty$, using strong $L^{2}$ convergence on the left hand side of (3.7), and pointwise a.e. convergence plus the dominated convergence theorem on the right hand side of (3.7), we see
$\int^{T}_{0}\langle x(t),\phi^{\prime}(t)\rangle dt=\int^{T}_{0}\langle f(x(t),u(t)),\phi(t)\rangle dt,\,\,\,\,\,\forall\phi\in C^{1}_{c}([0,T];{\mathbb{R}}^{6}),$
showing that $x$ is the solution of (2.2) corresponding to $u$. That is $\mathcal{S}(u)=x$, and $\mathcal{S}$ is weakly closed, which completes Step 1.
Step 2: $\mathcal{S}$ is weak-strong continuous.
Suppose $u^{(k)}\rightharpoonup u$ weakly in $L^{2}$. We need to prove that $\mathcal{S}(u^{(k)})\to\mathcal{S}(u)$ in $L^{2}.$ Once again, $\{\mathcal{S}(u^{(k)}\}$ is a uniformly bounded sequence in $(H^{1}[0,T])^{6}$, and so any subsequence $\{\mathcal{S}(u^{(k_{\ell})}\}$ is still uniformly bounded $(H^{1}[0,T])^{6}$. By the compact embedding $H^{1}[0,T]\hookrightarrow L^{2}[0,T]$, $\{\mathcal{S}(u^{(k_{\ell})}\}$ has a further subsequence $\{S(u^{(k_{\ell_{m}})}\}$ converging strongly (hence also weakly) to some function $x\in(L^{2}[0,T])^{6}$. By weak closure, we must have $\mathcal{S}(u)=x$, so we have proven that every subsequence of $\{\mathcal{S}(u^{(k)}\}$ has a further subsequence converging strongly to $\mathcal{S}(u)$ in $(L^{2}[0,T])^{6}$. This implies that $\mathcal{S}(u^{(k)})\to\mathcal{S}(u)$ strongly in $(L^{2}[0,T])^{6}$ by Urysohn’s subsequence principle [53, Ch. 2.1.17].
Step 3: The cost functional is lower-semicontinuous so (3.5) has a solution.
Because the control-to-state operator is well-defined, (3.5) is equivalent to the reduced problem
(3.8) $\min_{u\in\mathscr{U}_{\text{ad}}}\mathcal{J}(u)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=J(\mathcal{S}(u),u).$
Since $r,\ell$ are proper, convex, and lower-semicontinuous, they are bounded below, and since $u,\mathcal{S}(u)$ are bound bounded functions, this implies $\mathcal{J}(u)$ is bounded from below and thus has a finite infimum $\mathcal{J}_{0}$. Let $\{u^{(k)}\}$ be a minimizing sequence of controls, so that $\mathcal{J}(u^{(k)})\to\mathcal{J}_{0}.$ Since $\{u^{(k)}\}$ is a bounded sequence in the Hilbert space $(L^{2}[0,T])^{4}$, we can pass to a subsequence that converges weakly to some $u\in\mathscr{U}_{\text{ad}}$. By weak closure and weak-strong continuity, $\mathcal{S}(u^{(k)})\to x\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathcal{S}(u)$ strongly in $(L^{2}[0,T])^{6}.$ Since $r,\ell$ are proper, convex, and lower-semicontinuous, the maps
$x\mapsto\int^{T}_{0}r(x(t))dt,\,\,\,\,\,\,\,\,u\mapsto\int^{T}_{0}\ell(u(t))dt$
are weakly lower-semicontinuous, and thus
$\mathcal{J}_{0}\leq\mathcal{J}(u)=J(x,u)$ $=\int^{T}_{0}(r(x(t))+\ell(u(t)))dt$
$\leq\liminf_{k\to\infty}\int^{T}_{0}\Big((r([\mathcal{S}(u^{(k)})](t))+\ell(u^{(k)}(t))\Big)dt=\liminf_{k\to\infty}\mathcal{J}(u^{(k)})=J_{0}.$
Thus $u\in U_{\text{ad}}$ and the associated solution $x$ of (2.2) are a solution of (3.5). ∎
We prove one further regularity result for the control-to-state map. While this results was not necessary in the proof above, it will have bearing in section 4.
Lemma 3.2.
The control-to-state operator $\mathcal{S}:\mathscr{U}_{\text{ad}}\to(H^{1}[0,T])^{6}$ is Lipschitz continuous when both the source and target spaces are given the $L^{2}$ topology.
Proof.
Let $u,v\in\mathscr{U}_{\text{ad}}$ be two control maps, and let $x=\mathcal{S}(u)$ and $y=\mathcal{S}(v)$ be the associated solutions of (2.2). Then for any $t\in[0,T],$
$\left|x(t)-y(t)\right|_{2}\leq\int^{t}_{0}\left|f(x(s),u(s))-f(y(s),v(s))\right|_{2}ds.$
By proposition 2.1, solutions of (2.2) are uniformly bounded over choices of controls, and the controls themselves are uniformly bounded. Thus since $f(x,u)$ is smooth in both $x$ and $u$, it is, in particular, Lipschitz continuous, so
$\left|x(t)-y(t)\right|_{2}\leq L\int^{t}_{0}\Big(\left|x(s)-y(s)\right|_{2}+\left|u(s)-v(s)\right|_{2}\Big)ds.$
Thus Grönwall’s inequality gives
$\left|x(t)-y(t)\right|_{2}\leq Le^{T}\int^{t}_{0}\left|u(s)-v(s)\right|_{2}ds\leq Le^{T}\int^{T}_{0}\left|u(s)-v(s)\right|_{2}ds.$
Squaring this and invoking Jensen’s inequality yields
$\left|x(t)-y(t)\right|^{2}_{2}\leq L^{2}Te^{2T}\|u-v\|^{2}_{(L^{2}[0,T])^{4}}$
and integrating on $[0,T]$ shows that
$\|x-y\|^{2}_{(L^{2}[0,T])^{6}}\leq(LTe^{T})^{2}\|u-v\|^{2}_{(L^{2}[0,T])^{4}}$
proving the claim. ∎
Remark. A remark is in order regarding the specific choice of our control set $\mathscr{U}_{\text{ad}}$. If the dynamics are control affine, it is enough to assume $u\in(L^{2}[0,T])^{4}$. However, our dynamics are not control affine. Thus in Step 1 of the proof, some higher regularity is required to ensure a stronger notion of convergence of $u^{(k)}\to u$, and allow us to pass to limit in (3.7). The proof above would go through a bit more easily if we assume even more regularity–for example $u\in(H^{1}[0,T])^{4}$–but this is highly undesirable, because it would preclude the possibility of discontinuous control maps. Discontinuous controls arise naturally when enforcing $L^{1}$ control costs since these commonly lead to bang-bang control strategies (see [33, Ch. 17] or [52, 57]). The assumption that the controls have uniformly bounded total variation allows for the use of Helly’s selection theorem [25, Thm. 5.5] to pass to subsequence which converges pointwise a.e., while still maintaining strong closure of $\mathscr{U}_{\text{ad}}$ in the $L^{2}$ topology, and also allowing for discontinuous control maps.
@@
Given existence of optimal controls, a necessary condition for optimality is provided by the Pontryagin minimum principle (PMP). Using state variables $x=(S,I,R,S^{*},I^{*},R^{*})$, costate variables $p=(p_{S},p_{I},p_{R},p_{S^{*}},p_{I^{*}},p_{R^{*}})$, and control variables $u=(\alpha,\eta,\mu,\nu),$ we define the Hamiltonian
(3.9) $\begin{split}\mathcal{H}(x,p,u)&=\langle p,f(x,u)\rangle+r(x)+\ell(u),\end{split}$
@@
where $r,\ell$ are as in (3.1) and $f(x,u)$ is defined by the right-hand side of (2.2). Then the PMP states that if $x(\cdot),u(\cdot)$ are an optimal state and control pair for (3.5), then there is a corresponding costate trajectory $p(\cdot)$ such that